OpenStudy (anonymous):

In a mixture of concrete, there are two lb of cement mix for each pound of gravel. If the mixture contains a total of 120 lb of these two ingredients, how many pounds of gravel are there? I suck at word problems.

OpenStudy (anonymous):

2lbs of concrete / 1 lb of gravel * 120lbs total. 2 + 1 = 3 120 / 3 = 40 2 * 40 = 80 = concrete 1 *40 = 40 = gravel

OpenStudy (anonymous):

sorry, that should say 80 = cement

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

No, I'm confused @engstat